Use the indicated property to write an expression that is equivalent to the following expression.
(9 • 0) •17 (associative property of multiplication)

Answer by Edwin McCravy(8999)

Use the indicated property to write an expression that is equivalent to the following expression.
(9 • 0) • 17 (associative property of multiplication)
--------------------------------------------------------

Parentheses are used to ASSOCIATE two numbers, so in this

(9 • 0) • 17

The 9 and the 0 are ASSOCIATED by the parentheses around them.

Now let's move the parentheses so as to ASSOCIATE the 0 and the
17.  So we then have this:

9 • (0 • 17)

---------------------------------------------------------------

When we have three numbers to add or multiply, the fact that
it doesn't matter whether we associate the first two or the
last two, we'll get the same answer in the end -- is called the
ASSOCIATIVE property.

Example for the ASSOCIATIVE property of multiplication:

(2 • 5) • 7 = 10 • 7 = 70

 2 • (5 • 7) = 2 • 35 = 70

The work was different but we ended up with the same answer 70

Example for the ASSOCIATIVE property of addition:

(3 + 4) + 6 = 7 + 6 = 13

 3 + (4 + 6) = 3 + 10 = 13

Again the work was different but we ended up with the same answer 13

Edwin